INFRARED LASER SPECTROSCOPY OF METHANE-PARAHYDROGEN CLUSTER IN HELIUM DROPLETS

\underline{HIROMICHI HOSHINA}, MIKHAIL N. SLIPCHENKO, ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA.

The clusters of methane with para-hydrogen molecules (CH_{4}-(p-H_{2})_{n} n=1∼30) were formed in He droplets and studied via laser depletion spectroscopy. The rotationally resolved ν_{3} transition of the cluster was observed with changing the number of hydrogen molecules. Large spectral change was seen around n=12 and 30, which suggests the completion of the first and second H_{2} solvation shells of the CH_{4} molecule. The aggregate state of the H_{2} cluster is discussed.